TheSmartOne (thesmartone):

Evaluate the limit as x approaches 0 of (1 - x^(sin(x)))/(x*log(x)) aka: \(\frac{1-x^{sin(x)}}{x*log(x)}\) Please include steps/explanation.

TheSmartOne (thesmartone):

@ganeshie8 @Michele_Laino

OpenStudy (er.mohd.amir):

use l'Hospital rrule

TheSmartOne (thesmartone):

\(\LARGE \frac{1-x^{sin(x)}}{x*log(x)}\)

TheSmartOne (thesmartone):

\(\large \lim_{x \to 0^{+}} \frac{1- x^{x} }{ \log( x^x) } =_{LH} \lim_{x \to 0^{+}} \frac{0 -x^x( 1 + \log (x)) }{1 + \log (x) } \\ = \large \lim_{x \to 0^{+}} (-x^x) = \large - \lim_{x \to 0^{+}} (x^x) = -1\)

TheSmartOne (thesmartone):

@Er.Mohd.AMIR correct?

TheSmartOne (thesmartone):

@sleepyhead314 I know you know the answer ;)

OpenStudy (jango_in_dtown):

OK I can solve the problem

OpenStudy (sleepyhead314):

cannot directly use l'hopital

TheSmartOne (thesmartone):

yea, before that I had this; \(\large \lim_{x \to 0^{+}} \frac{1- x^{\sin x} }{x \log x } \\~\\ \large = \lim_{x \to 0^{+}} \frac{1- x^{\sin x} }{ \log( x^x) } \\~\\ \large = \lim_{x \to 0^{+}} \frac{1- x^{x} }{ \log( x^x) } ~~ \normalsize{\text{ substituting x for sin x } } \\~\\ \large = \frac{\lim_{x \to 0^{+}} (1) - \lim_{x \to 0^{+}} \left( x^{x}\right) }{ \log( \lim_{x \to 0^{+}}x^x) } = \frac{1-1}{\log(1)} = \frac{0}{0}\)

OpenStudy (sleepyhead314):

I know that you know that you have absolutely no idea what that means

TheSmartOne (thesmartone):

or do I (;

OpenStudy (er.mohd.amir):

wrong diff of numerator.

OpenStudy (jango_in_dtown):

check my solution

TheSmartOne (thesmartone):

@hartnn :)

OpenStudy (anonymous):

calculus problem lol

OpenStudy (just_one_last_goodbye):

@TheSmartOne did you tag Hartnn for help or for CoC enforcement? Cause I can help with the question.

TheSmartOne (thesmartone):

for help :)

OpenStudy (jango_in_dtown):

@ganeshie8 does my solution contains error? do let me know

TheSmartOne (thesmartone):

if you can help @just_one_last_goodbye , go for it :)

OpenStudy (just_one_last_goodbye):

working it out on my paper :)

TheSmartOne (thesmartone):

sure, tell me what you get :)